# SelfElicit: Your Language Model Secretly Knows Where is the Relevant Evidence

### v1

@@ B.1 Computational efficiency
Due to space constraints, we report only the average running time per sample across all datasets in Table 1. Table 5 provides a detailed breakdown of efficiency results for each dataset, including the average inference time per sample (i.e., context-question pair) and the total number of tokens generated during inference. We observe that: (i) While FullElicit is highly efficient as a naive method, it barely helps with the QA performance, as shown in Table 1. (ii) PromptElicit, based on generative evidence extraction, requires the model to generate a large volume of extra tokens during inference, up to 852% more than direct QA, resulting in considerably slower inference times. (iii) As discussed before, the response to Cot is not consistent for different LMs. Generally, for Llama and Mistral models, Cot have little impact on the QA performance, and inference time as well. For Qwen-series models, Cot obtains slightly better EM by generating the intermediate reasoning steps before answering, but this also leads to a great drop in F1 score and significantly higher inference time (e.g., +72.0% for Qwen2.5-7B) due to the generation of intermediate step tokens. (iv) SelfElicit, by helping the model focus on contextual evidence, generates shorter but more accurate answers. Additionally, as SelfElicit does not involve generative evidence extraction, it remains computationally efficient, with only 3-5% of the inference time overhead compared to PromptElicit.
@@ B.2 More layer-wise attention visualization
B.3 Sentence-level versus token-level eliciting
@@
B.4 Examples with full context
Finally, we present the full context of examples from Table 2 to illustrate SelfElicit’s ability to identify supporting facts/evidence within noisy long contexts. The results are shown in Tables 7 and 8. These examples are from HotpotQA’s distractor setting, where the context passage contains distracting information retrieved from Wikipedia using the question as a query Yang et al. (2018). Note that these distractors are not completely irrelevant random noise that can be easily filtered out by retrieval systems. Instead, they appear related to the question but do not actually support answering it, serving as "hard negatives." This scenario is quite common in practice and significantly impacts the performance of retrieval-augmented generation Wu et al. (2024); Cuconasu et al. (2024); He et al. (2024); Liu et al. (2024b). As shown in Tables 7 and 8, the LM struggles to effectively use contextual evidence to provide correct answers under such a situation. Despite this, SelfElicit accurately highlights the critical evidence within the noisy long context, helping the model focus on the most relevant facts and thus arrive at the correct answer.
Appendix C Additional Discussions
C.1 Potential Risks
Generative AI tools such as language models have an increasing impact on our daily lives in the era of big data and AI Yan et al. (2024); Xu et al. ; Ban et al. (2021); Lin et al. (2024a, b); Guo et al. (2023), such as finance Chan et al. (2024); Liu et al. (2024d) and healthcare Ye et al. (2023); Liu et al. (2024c), especially with the recent trends of foundation models Zheng et al. (2024a); Fu et al. (2024); Zheng et al. (2024b); Li et al. (2024). This study focuses on enhancing language models’ ability to effectively utilize information from contextual documents. However, retrieved documents from the internet may contain unethical or discriminatory content, which the model might read and incorporate into its outputs. While addressing ethical concerns in Retrieval-Augmented Generation (RAG) or fact-based question-answering tasks Liu et al. (2024a, 2022, 2021a) is beyond the scope of this work, such issues can typically be mitigated by using detectors to filter harmful information from the context documents provided to the language model, or other general techniques such as ensemble multi-model answers Liu et al. (2021b, 2020a, 2020b) or watermarking Chen et al. (2024a) AI-generated contents.
C.2 Usage of Artifacts and AI Assistants

### v2

@@ B.1 Computational Efficiency
Due to space constraints, we report only the average running time per sample across all datasets in Table 1. Table 6 provides a detailed breakdown of efficiency results for each dataset, including the average inference time per sample (i.e., context-question pair) and the total number of tokens generated during inference. Note that Mistral-12B generates shorter answers (7.34 tokens on average) for each sample compared to Mistral-7B (24.06 tokens), thus showing a shorter inference time for each sample as it generates fewer tokens. We observe that: (i) While FullElicit is highly efficient as a naive method, it barely helps with the QA performance, as shown in Table 1. (ii) PromptElicit, based on generative evidence extraction, requires the model to generate a large volume of extra tokens during inference, up to 852% more than direct QA, resulting in considerably slower inference times. (iii) As discussed before, the response to Cot is not consistent for different LMs. Generally, for Llama and Mistral models, Cot have little impact on the QA performance, and inference time as well. For Qwen-series models, Cot obtains slightly better EM by generating the intermediate reasoning steps before answering, but this also leads to a great drop in F1 score and significantly higher inference time (e.g., +72.0% for Qwen2.5-7B) due to the generation of intermediate step tokens. (iv) SelfElicit, by helping the model focus on contextual evidence, generates shorter but more accurate answers. Additionally, as SelfElicit does not involve generative evidence extraction, it remains computationally efficient, with only 3-5% of the inference time overhead compared to PromptElicit.
@@ B.2 More Layer-wise Attention Visualization
B.3 Impact of Evidence Highlighting on Model Attention
To assess the effectiveness of SelfElicit in directing model attention toward relevant context, we analyzed the average attention score per token for the selected evidence sentences before and after highlighting. We conducted this analysis across four QA datasets using Llama-3.1-8B, averaging attention scores over all layers and heads. As shown in Table 7, we observe a consistent increase in attention scores for the highlighted evidence across all datasets. This demonstrates that the highlighting strategy effectively shifts model focus toward critical contextual information, reinforcing its ability to retrieve and utilize evidence more effectively.
Table 7: Effect of highlighting on model attention, measured as average attention per token (APT) for selected evidence sentences.
APT (×$\times 10^{-3}$) HotpotQA NewsQA TQA NQ
Before Highlighting 0.4362 0.3465 0.2090 0.9631
After Highlighting 0.6351 0.5783 0.3451 1.1917
After/Before Ratio 145.61% 166.91% 165.10% 123.73%
B.4 Sentence-level versus Token-level Eliciting
@@
B.5 Alternative Highlighting Strategies
In our primary experiments, we utilized in-context highlighting, where relevant evidence is marked directly within the original passage. This approach aligns with the core insight of SelfElicit —leveraging the LM’s inherent ability to identify relevant evidence—while maintaining the semantic continuity of the input. However, alternative methods of evidence integration could potentially influence the model’s ability to utilize highlighted information effectively. To explore this, we conducted additional experiments using non-in-context highlighting, where all extracted evidence sentences were aggregated and appended either to the beginning or end of the original context. We evaluated these variations using Llama-3.1-8B across four datasets and reported F1 scores in Table 9.
Table 9: Comparison of different highlighting strategies.
Highlighting HotpotQA NewsQA TQA NQ
- 57.7 56.2 66.1 61.6
In-context 69.5 60.8 72.7 67.8
Attach at Begin 69.8 59.8 69.3 67.8
Attach at End 69.0 58.8 67.0 65.6
The results indicate that in-context highlighting generally yields the best performance, likely due to its preservation of the original passage’s semantic flow, allowing the model to process highlighted evidence within its natural context. Appending evidence to the beginning or end also improves performance over the base model but is slightly less effective, particularly for multi-hop tasks like HotpotQA, where contextual relationships between facts are crucial.
We emphasize that SelfElicit adopts in-context highlighting primarily for its simplicity and efficiency. However, our framework is flexible, and alternative highlighting strategies can be seamlessly incorporated. Exploring more effective context augmentation techniques remains an intriguing direction for future work.
B.6 Potentially Uninformative Context
To further analyze SelfElicit’s robustness in handling cases where the provided context may lack relevant evidence, we constructed an additional QA task based on HotpotQA. This experiment assesses whether SelfElicit can correctly reject answers when the context does not support a response. Specifically, we randomly selected 50% of the samples as unanswerable, replacing their original context with unrelated passages from other samples. The correct response in these cases was set to: "I cannot answer based on the given context." We explicitly instructed the LLM to generate this response whenever relevant information was absent. In addition to standard EM and F1 scores, we introduced Rejection Accuracy, which measures the model’s ability to correctly identify and reject uninformative contexts. The results on Llama-3.1-8B are shown in Table 10.
Table 10: Results on potentially unanswerable QA task.
Method F1 EM RejectionAccuracy
Base 78.48 79.58 91.25
SelfElicit 82.65 81.93 92.86
These findings, along with Table 1 in the main paper, demonstrate that SelfElicit does not degrade—rather, it slightly enhances—the model’s ability to recognize when the provided context lacks sufficient information for answering. At the same time, SelfElicit significantly improves the model’s ability to utilize relevant information when it is available by highlighting the most crucial evidence in relation to the user query.
B.7 Highlighting versus Filtering Strategies
In this section, we compare SelfElicit to the Filco method which filters context using a fine-tuned Llama-2 model Wang et al. (2023b). Additionally, we explore the impact of highlighting versus filtering evidence. To analyze the impact of different evidence extraction and highlighting strategies, we conduct experiments on four SelfElicit/Filco variants: (i) SelfElicit-highlight: The standard SelfElicit approach, where identified evidence sentences within the context are explicitly highlighted. (ii) SelfElicit-filter: Instead of highlighting, this method extracts the identified evidence sentences and concatenates them to form a new, filtered context passage. (iii) Filco-filter: The standard Filco approach, which employs a fine-tuned Llama-2 model to generate a filtered version of the context for answering questions. (iv) Filco-highlight: This variant highlights the evidence sentences identified by Filco within the original context, rather than filtering them out.
This experimental setup enables a direct comparison between highlighting vs. filtering, as well as SelfElicit vs. Filco, in their ability to enhance LMs’ utilization of contextual information. We use Llama-3.1-8B-Instruct as the base LM and evaluate performance across four QA datasets. The F1 scores are presented in Table 11.
Table 11: Comparison of SelfElicit and Filco variants on four QA datasets, evaluated using F1 scores. SelfElicit-highlight achieves the highest performance across all datasets.
Method HotpotQA NewsQA TQA NQ
Base 57.7 56.2 66.1 61.6
SelfElicit-highlight 69.5 60.8 72.7 67.8
SelfElicit-filter 62.4 55.6 66.5 65.1
Filco-highlight 67.8 57.2 68.9 65.3
Filco-filter 62.6 54.7 66.3 64.5
The key findings are: (i) SelfElicit consistently achieves the highest performance across all datasets, outperforming not only its filtering-based counterpart but also Filco, a more complex, fine-tuned approach. This highlights the effectiveness of inference-time evidence elicitation over training-based filtering. (ii) Highlighting-based methods outperform their filtering-based counterparts for both SelfElicit and Filco. Filtering removes context that may contain auxiliary or indirect supporting information, such as intermediate reasoning steps, leading to information loss. (iii) Filtering-based methods struggle particularly on NewsQA, which consists of news articles as context. Unlike structured knowledge sources such as Wikipedia, news articles tend to be semantically cohesive—making it harder to extract self-contained and sufficiently informative evidence sentences. This may explain why Filco was originally evaluated on Wikipedia-based datasets, where evidence extraction is more straightforward.
These findings reinforce that highlighting key evidence within the original context is more effective than filtering, as it preserves both primary evidence and surrounding contextual cues that may contribute to reasoning. Moreover, SelfElicit, despite being a lightweight inference-time method, surpasses Filco, which relies on additional model fine-tuning—demonstrating the advantage of leveraging pretrained LMs’ inherent ability to locate and utilize relevant evidence.
B.8 Examples with full context
Finally, we present the full context of examples from Table 2 to illustrate SelfElicit’s ability to identify supporting facts/evidence within noisy long contexts. The results are shown in Tables 12 and 13. These examples are from HotpotQA’s distractor setting, where the context passage contains distracting information retrieved from Wikipedia using the question as a query Yang et al. (2018). Note that these distractors are not completely irrelevant random noise that can be easily filtered out by retrieval systems. Instead, they appear related to the question but do not actually support answering it, serving as "hard negatives." This scenario is quite common in practice and significantly impacts the performance of retrieval-augmented generation Wu et al. (2024); Cuconasu et al. (2024). As shown in Tables 12 and 13, the LM struggles to effectively use contextual evidence to provide correct answers under such a situation. Despite this, SelfElicit accurately highlights the critical evidence within the noisy long context, helping the model focus on the most relevant facts and thus arrive at the correct answer.
Appendix C Additional Discussions
C.1 Failure Case Analysis and Potential Improvements
While SelfElicit significantly enhances LM’s ability to identify and utilize relevant evidence, there are cases where it fails to aid the model in generating correct answers. Through manual analysis, we found that incomplete evidence highlighting is a primary cause of failure. Consider the following example where Llama-3.1-8B, both with and without SelfElicit, gives a negative response.
An example case with $\alpha=0.5$ Context: Alexander S. Wadsworth Commodore Alexander Scammel Wadsworth (1790–April 5, 1851) was an officer of the United States Navy. His more than 40 years of active duty included service in the War of 1812. War of 1812 The War of 1812 (1812–1815) was a conflict fought between the United States, the United Kingdom, and their respective allies. Historians in Britain often see it as a minor theater of the Napoleonic Wars; in the United States and Canada, it is seen as a war in its own right … Question: Alexander S. Wadsworth served during the war that was fought between which two countries? Base Model: I cannot answer based on the given context.
SelfElicit:
I cannot answer based on the given context.
We can see that although SelfElicit successfully highlighted the sentence: "The War of 1812 (1812–1815) was a conflict fought between the United States, the United Kingdom, and their respective allies." as important evidence, this alone was insufficient for the model to infer that Alexander S. Wadsworth served in this war. We observed that lowering $\alpha$ can mitigate such failures by allowing SelfElicit to highlight more comprehensive evidence. Reducing $\alpha$ to 0.1 enabled the method to additionally highlight the missing link—“Commodore Alexander Scammel Wadsworth (1790–April 5, 1851) was an officer of the United States Navy. His more than 40 years of active duty included service in the War of 1812.”, and lead to the correct answer.
An example case with $\alpha=0.1$ Context: Alexander S. Wadsworth Commodore Alexander Scammel Wadsworth (1790–April 5, 1851) was an officer of the United States Navy. His more than 40 years of active duty included service in the War of 1812. War of 1812 The War of 1812 (1812–1815) was a conflict fought between the United States, the United Kingdom, and their respective allies. Historians in Britain often see it as a minor theater of the Napoleonic Wars; in the United States and Canada, it is seen as a war in its own right … Question: Alexander S. Wadsworth served during the war that was fought between which two countries? Base Model: I cannot answer based on the given context.
SelfElicit:
United States, the United Kingdom, and their respective allies.
This suggests that a fixed threshold may not be optimal across different questions and contexts, and SelfElicit could potentially benefit from a more adaptive thresholding mechanism: Instead of using a fixed $\alpha$, a dynamic threshold that adjusts based on the complexity of the input context or question could provide more flexible evidence selection. Nevertheless, as shown and discussed in Figure 4, SelfElicit remains robust to the choice of $\alpha$, with a broad range of values yielding consistently strong improvements. Even with static thresholding, SelfElicit provides stable and significant gains in LM performance across QA tasks.
C.2 Understanding Layer-Wise Evidence Utilization in Language Models
Our analysis reveals that while intermediate layers of language models effectively emphasize relevant evidence within the context, this emphasis does not necessarily translate into correct final responses. As shown in Figure 3 and 5, attention to key evidence tends to diminish from mid-layers to the final layers, indicating that the final layers do not prioritize contextual evidence as strongly as mid-layers do. This suggests that although models inherently recognize relevant information at intermediate layers, their final token selection during decoding is influenced by additional factors that may override this evidence, leading to incorrect responses.
This phenomenon aligns with recent findings in closed-book QA and controlled generation settings. Prior work has shown that factually correct tokens often receive higher decoding probabilities in intermediate layers, while incorrect tokens can maintain consistently high probabilities across all layers and still be selected as the final output Chuang et al. (2024). Similarly, studies on the "overthinking" behavior of LMs indicate that removing the last few layers can sometimes improve factual accuracy, implying that deeper layers may introduce unnecessary complexity or spurious reasoning Halawi et al. (2024). Our findings provide further empirical support for these observations, explaining why models can correctly identify relevant evidence yet still produce incorrect answers.
C.3 Potential Risks
Generative AI tools such as language models have an increasing impact on our daily lives in the era of big data and AI Yan et al. (2021, 2022, 2023a, 2023b, 2024a, 2024b, 2024c); Xu et al. ; Xu et al. (2024); Ban et al. (2021); Lin et al. (2024a, b); Guo et al. (2023); Wei et al. (2025c, 2024b); Qiu et al. (2024c, b, a, 2023, 2022); Zeng et al. (2023a, b, 2024a, 2024b); Qiu and Tong (2024); Wei et al. (2025a); Ai et al. (2023); Yoo et al. (2025a, b, 2024); He et al. (2023); Bao et al. (2023); Wei et al. (2023, 2024a); Wei and He (2022); Wei et al. (2022b, 2020, 2021); Li et al. (2025), such as finance Chan et al. (2024); Liu et al. (2024c) and healthcare Ye et al. (2023); Liu et al. (2024b), especially with the recent trends of foundation models Zheng et al. (2024a); Fu et al. (2024); Zheng et al. (2024b); Li et al. (2024b); Zhang et al. (2025); Ai et al. (2025). This study focuses on enhancing language models’ ability to effectively utilize information from contextual documents. However, retrieved documents from the internet may contain unethical or discriminatory content, which the model might read and incorporate into its outputs. While addressing ethical concerns in Retrieval-Augmented Generation (RAG) or fact-based question-answering tasks Liu et al. (2024a, 2022, 2021a) is beyond the scope of this work, such issues can typically be mitigated by using detectors to filter harmful information from the context documents provided to the language model, or other general techniques such as ensemble multi-model answers Liu et al. (2021b, 2020a, 2020b) or watermarking Chen et al. (2024a) AI-generated contents.
C.4 Usage of Artifacts and AI Assistants